OpenStudy (anonymous):

A university bookstore recently sold a wirebound graph-paper notebook for $1.10, and a college-ruled notebook for $3.16. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $102.38. How many of each type were sold?

hero (hero):

w + c = 50 1.10w + 3.16c = 102.38 Solve the system

OpenStudy (anonymous):

I am confused

hero (hero):

Confused about what?

OpenStudy (anonymous):

about how to do it

hero (hero):

Basically you write both equations in terms of w, set w = w, cross multiply, solve for c, then solve for w: w + c = 50 1.10w + 3.16c = 102.38 w = 50 - c w = (102.38 - 3.16c)/1.10 w = w 1.10(50 - c) = 102.38 - 3.16c 55 - 1.10c = 102.38 - 3.16c -1.10c + 3.16c = 102.38 - 55 2.06c = 47.38 c = 47.38/2.06 c = 23 Do you think you can find w?

OpenStudy (anonymous):

w=50-23=27

hero (hero):

correct

OpenStudy (anonymous):

ok so is that the answer

hero (hero):

What do you think?

hero (hero):

If you don't understand my steps, you should consider consulting a tutor.

OpenStudy (anonymous):

well do I need to work this out as well w = (102.38 - 3.16c)/1.10

hero (hero):

You only needed to solve for c and w. You have both values now. c = 23, w = 27. The only thing you need to do is study and understand the steps I've given you.

OpenStudy (anonymous):

oh ok